The following information is from Pramacy Inc.’s financial statements:

Sales (all credit)

$245,000

Total assets turnover

0.80 times

Total debt to total assets

30.83%

Fixed asset turnover

1.1 times

Current ratio

4.04 times

Average collection period

29.68 days

Inventory turnover

7.42 times

Part A

Using the above listed ratios and data, compute the balance of the following accounts. Assume all sales are on credit and a 360-day year. Round to the nearest dollar. [Hint: use ratio formulas to derive the requested values. The book presents 13 ratio formulas on pages 60 - 64]. Please show your work.

64)

a) Total asset

b) Total debt

c) Total equity

d) Fixed assets

e) Current assets

f) Average daily credit sales

g) Accounts Receivable

h) Inventory

i) Current liabilities
Answer #1

As per rules I am answering the first 4 subparts of the question

a) Total asset = Sales/ TAT

= 245000/0.8

= 306250

b) Total debt = 30.83%*Total assets

= 30.83%*306250

= 94416.88

c) Total equity = 306250* (100%-30.83%)

= 211833.1

d) Fixed assets = Sales*Fixed asset turnover

= 245000/1.1

= 222727.27